OpenStudy (anonymous):

Are 3 times x plus 33 all over x plus 11 and 3 equivalent? Using complete sentences explain why or why not.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

is the question "does \(\frac{3x+33}{x+11}=3\)? "

OpenStudy (ashleyisakitty):

(3x+33)/{(x+11)*3} {3*(x + 11)}{(x+ 11)*3} (3/3)*(x + 11)/(x+11) 1*1 1

OpenStudy (skullpatrol):

Think about when x = -11 @ahoward79

OpenStudy (skullpatrol):

@ahoward79 are you still there?

OpenStudy (anonymous):

Im so confused..

OpenStudy (skullpatrol):

About what?

OpenStudy (ashleyisakitty):

Ahoward, i solved out the equation for you. Tell me what you have to say about the ending of it? Does it seem equivalent?

OpenStudy (skullpatrol):

You know that division by 0 is undefined, right?

OpenStudy (anonymous):

No it doesnt, but I dont really understand how you solved it.

OpenStudy (anonymous):

@Ashleyisakitty it is not 1, but rather 3

OpenStudy (anonymous):

\[\frac{3x+33}{x+11}=\frac{3(x+11)}{x+11)}=\frac{3\cancel{(x+11)}}{\cancel{x+11}}=3\] which will be true for all \(x\) unless \(x=-11\) in which case the original expression is \[\frac{0}{0}\] which is undefined